OpenStudy (lukecrayonz):

Find the area under the given curve over the indicated interval y=6 [-2,2]

OpenStudy (anonymous):

Thats essentially a square. But you can also take the integral over that time interval and get the same answer

OpenStudy (anonymous):

i would say it was a rectangle

OpenStudy (anonymous):



OpenStudy (lukecrayonz):

24?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

\[\int\limits_{-2}^{2} 6 dx = [6x]_{-2}^{2}= 12-(-12)=24\]